if x varies inversely as v, and x=14 when v=2, find x when v=4.
Answer by nerdybill(7384)

if x varies inversely as v, and x=14 when v=2, find x when v=4.
.
When you see "x varies inversely as v"
You're suppose to automatically think:
x = k/v
where k is a constant
.
To find k, we use "x=14 when v=2":
x = k/v
14 = k/2
28 = k
.
We now have this as the relation:
x = 28/v
.
Finally, we can answer: "find x when v=4"
x = 28/v
x = 28/4
x = 7